OpenStudy (yuii):

Select all numbers that are irrational. √144 √8 16π 0.45

OpenStudy (knawk):

Do you know what irrational numbers are?

OpenStudy (yuii):

Most of them.

OpenStudy (knawk):

Anything with pi, and under square root that cant be simplified to one whole number, and also fractions and decimals.

OpenStudy (mysticfury):

an irrational number is a decimal that does not repeat

OpenStudy (mysticfury):

and goes on for a long time, its basically anything that cant be put as a fraction

OpenStudy (knawk):

there's only 1 number that isn't irrational, which do u think it is?

OpenStudy (yuii):

C.

OpenStudy (knawk):

pi = 3.14 16*3.14 would give us a decimal, so it's irrational. A is the correct answer. it simplifies to 12

OpenStudy (eliesaab):

The irrationals are \[ \sqrt 8\\ 16 \pi \]

OpenStudy (knawk):

I'm sorry, i meant A is the only one that isn't irrational. All the rest are your correct answers.

OpenStudy (eliesaab):

\[\pi\] is not 3.14 3.14 is an approximation of \(\pi\)

OpenStudy (yuii):

Now I'm confused. Is all of them but A irrational or just B and C?

OpenStudy (eliesaab):

only the two I mentioned are irrationals \[ \sqrt {144}=12\\ .45=\frac{45}{100} \]

OpenStudy (knawk):

@eliesaab any number times pi is the same as times 3.14.

OpenStudy (knawk):

anyway, numbers times pi will give you a decimal, so pi is irrational.

OpenStudy (eliesaab):

any number rational number times \(\pi\) is irrationals

OpenStudy (knawk):

@Yuii Sorry for the confusion, your correct answers are B,C,D

OpenStudy (yuii):

Are you sure? This is my last attempt on this assignment.

OpenStudy (knawk):

I'm positive, B,C, and D are IRRATIONAL. A is RATIONAL. Double check from the question please.

OpenStudy (eliesaab):

Here is pi with 200 decimal places \[ \pi =3.14159265358979323846264338327950288419716939937510582097494459230781 64062862089986280348253421170679821480865132823066470938446095505822317253594081284811174502841027019385211055596446229489549303819644288109757 \]

OpenStudy (knawk):

@eliesaab Lol, you round it, if you round it it would end up being 3.142 the 2 isn't much of a value since it's at the 1000th. Ask your math teacher if you don't believe me.

OpenStudy (eliesaab):

Your answer should be B and C only. These are the irrational numbers

OpenStudy (knawk):

Eliesaab, D is a fraction...

OpenStudy (knawk):

decimal*

OpenStudy (eliesaab):

An irrational number is by definition a non repeating decimal like \(\pi\) and \(\sqrt 2\) etc

OpenStudy (knawk):

lol

OpenStudy (eliesaab):

You do not have to agree with me. That is your choice of getting the answer wrong

OpenStudy (knawk):

Eliesaab can you recap the correct answers?

OpenStudy (eliesaab):

B and C are the correct answer. Period

OpenStudy (knawk):

Ok, hope that's the correct answer.

OpenStudy (eliesaab):

It is.

OpenStudy (knawk):

@eliesaab but can you please read this?

OpenStudy (eliesaab):

I read it and so what

OpenStudy (knawk):

decimals do not exist on the *number line*

OpenStudy (eliesaab):

the number line has all real numbers, rational and irrationals, A rational number can be written as the quotient of two integers. An irrational number cannot be written as a quotient of two integers,

OpenStudy (knawk):

Proof says decimals do not exist on the number line. I'm not gonna debate further. Perhaps you take the say in this answer? :)

OpenStudy (eliesaab):

http://www.dictionary.com/browse/number-line

OpenStudy (knawk):

lol? that doesn't say anything about decimals being rational.